OpenStudy (anonymous):

A bucket of apples weighs 6 1/2 pounds. If the apples are going to be divided up evenly among four friends, how many pounds do they each receive?

OpenStudy (anonymous):

okay.. are there options?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay do you know what 6 1/2 is as a decimal?

OpenStudy (anonymous):

6.5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a. 1 1/2 b. 1 5/8 c. 26 d. 1 3/4

OpenStudy (anonymous):

It's C

OpenStudy (anonymous):

i think it may be B. but why do you think that i want to see if i am wrong?

OpenStudy (anonymous):

of course they want it evenly between 4 people yes but i m thinking that you divide 4 by 6.5 and get 1.625 so to me it would be B. i am unsure

OpenStudy (anonymous):

i mean 6.5 by 4

OpenStudy (anonymous):

then when you divide 5/ 8 you get .625

OpenStudy (anonymous):

\[6 \frac{ 1 }{ 2}=\frac{ 4 }{ 1 }= \frac{ 52 }{ 2 }= 26?\]

OpenStudy (anonymous):

\[6 \frac{ 1 }{ 2 }= \frac{ 13 }{ 2 }= \frac{ 4 }{ 1 }=\frac{ 52 }{ 2}= 26?\]

OpenStudy (anonymous):

you know you cannot simplify 13 bc it is a prime right?

OpenStudy (anonymous):

u have your book right?? what does it say?

OpenStudy (anonymous):

\[\frac{ 13 }{ 2}{\times } \frac{ 4 }{ 1 }= \frac{ 52 }{ 2 }= 26\]

OpenStudy (anonymous):

Ohhhh I get it

OpenStudy (anonymous):

yea..

OpenStudy (anonymous):

so r u right or am i?

OpenStudy (anonymous):

Bye!!

OpenStudy (anonymous):

yes thank you byee